Circular No. 8622

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2000 QJ_{46} (LINEAR) An apparently asteroidal discovery by the LINEAR project (discovery

observation below from MPS 18126) has been reported now by M. Solontoi and A. West (Astronomy Department, University of Washington; with assistance by D. Schlelgel) as having a cometary appearance (with a coma and a very faint, short tail) on Sloan Digital Sky Survey 2.5-m telescope (Apache Point, NM) frames from 2000 Sept. 3 and 4; the unpublished SDSS astrometry is added below; the comet showed the following SDSS magnitudes on Sept. 3.27 UT: u = 19.6, g = 18.0, r = 17.4, i = 17.2, z = 17.1.

2000 UT α_{2000} δ_{2000} Mag. Observer
Aug. 24.27007 22^{h}34^{m}38.^{s}81 −8^{o}24′11′′.5 18.5 LINEAR
Sept. 3.26725 22 30 05.10 −9 24 26.4  SDSS
4.38530 22 29 33.65 −9 31 21.3  _{”}

The following orbital elements are from MPO 13685:

Epoch = 2001 Apr. 1.0 TT

T = 2000 Dec. 10.3467 TT  ω = 222.^{o}9660
e = 0.673150  Ω = 158.2344 }2000.0
q = 1.933644 AU  i = 4.3757

a = 5.915996 AU n^{o} = 0.0684955 P = 14.39 years

SUPERNOVA 2005hf R. Quimby, P. Hoeflich, S. J. Kannappan, and J. C. Wheeler, Uni-

versity of Texas; and C. Gerardy, Imperial College, report the discovery of a supernova (at mag ∼ 16.7) in unfiltered CCD images taken on Oct. 25.12 UT using the 0.45-m ROTSE-IIIb telescope at the McDonald Observatory. The new object, found by subtracting a co-addition of images taken between July 16 and Sept. 12 (limiting mag ∼ 18.9), is located at α = 1^{h}27^{m}05^{s}.97, δ = +19^{o}07′00′′.5 (equinox 2000.0; uncertainty ± 0′′.5), which is 1′′.9 west and 1′′.5 north of the apparent host galaxy. A spectrogram (range 420–890 nm) of SN 2005hf, obtained on Oct. 25.35 with the 9.2-m Hobby-Eberly Telescope (+ Marcario Low-Resolution Spectrograph) by J. Caldwell and E. Terrazas, shows it to be a type-Ia supernova near maximum light. In addition to Si II 635.5-nm, S II 545.4- and 564.0-nm, and weak Ca II infrared absorption features, fairly strong Na I absorption is detected that is consistent with the NED recession velocity of 12924 km/s (from Maurogordato et al. 1997, A.Ap. Suppl. 123, 411).

2005 October 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green